# Eleanor Wiske Dillon Sonia Jaffe Microsoft Research Microsoft Research Nicole Immorlica Christopher T. Stanton Microsoft Research Harvard Business School, NBER, and CEPR

### v1

@@ Abstract
We present evidence on how generative AI changes the work patterns of knowledge workers using data from a 6-month-long, cross-industry, randomized field experiment. Half of the 6,000 workers in the study received access to a generative AI tool integrated into the applications they already used for emails, document creation, and meetings. We find that access to the AI tool during the first year of its release primarily impacted behaviors that could be changed independently and not behaviors that required coordination to change: workers who used the tool spent 3 fewer hours, or 25% less time on email each week (intent to treat estimate is 1.4 hours) and seemed to complete documents moderately faster, but did not significantly change time spent in meetings.
@@
Recent advances in generative AI have opened new possibilities for using AI to assist with or automate a variety of tasks. Early studies have already shown that these technologies can substantially increase workers’ productivity in targeted tasks (Brynjolfsson, Li and Raymond (2025), Peng et al. (2023), Noy and Zhang (2023), Dell’Acqua et al. (2023)). Over time, these advances have the potential to shift the nature of jobs (Eloundou et al., 2024) and perhaps the macroeconomy (Autor (2024), Acemoglu (2025)). However, the transition from the creation of new technologies to effective, widespread adoption in the workplace takes time (Brynjolfsson, Rock and Syverson, 2021). In lab experiments, participants are asked to focus on completing a single task and directed to use AI tools that are well-suited to that task. Outside the lab, workers must learn on their own which parts of their job benefit from the use of new tools and gradually change their habits of work. Larger productivity gains often require organizational changes to redesign processes and shift responsibilities across workers (Bresnahan, Brynjolfsson and Hitt, 2002).
@@
In this paper, we present early evidence on how generative AI tools are changing the way knowledge workers do their jobs. Between September 2023 and October 2024, Microsoft partnered with 56 large firms to run a cross-industry field experiment to measure how access to an integrated generative AI tool, Microsoft 365 Copilot (hereafter, Copilot), changed work patterns. Each firm recruited workers for the study; workers were then randomly assigned to either receive access to this new generative AI tool or continue working with their current technologies. Participants in the study span many occupations, but all participating workers do some kind of work that involves heavy use of the Microsoft Office tools into which Copilot was integrated (e.g., emailing, video meetings, chat, document creation). Firms maintained the random assignment for six months, giving treated workers sufficient time to explore the new tool and integrate it into their daily work.
@@
This study took place during the early rollout of Copilot, when firms were restricted to purchasing a small number of licenses [^1] Each company was able to purchase 300 licenses. In all cases, this represents only a small fraction of their total employees who used Microsoft’s Office suite.. This particular generative AI tool was new for all workers, and the modal worker in the study had no close coworkers with access to the tool. Firms varied in how much training in the new tool they provided to treated workers. These workers were then free to use or not use the tool during their regular work. Treated and control workers may have also used other generative AI tools during the study, for example ChatGPT had been publicly available for nearly a year by October 2023, but access to Copilot could make it easier for workers to use generative AI at work. Copilot integrates AI into applications these workers were already using regularly in their jobs, allowing some new uses, for example summarizing emails or searching for information within work content, and reducing the effort cost of using generative AI to create content within documents and emails.
@@
We used highly detailed telemetry data collected by Microsoft products to track the behavior of the over 6,000 studied workers during the experiment and for several months before the introduction of the new tool. These telemetry measures capture the actions of workers, for example time spent reading emails or attending video meetings or time from creation to completion of documents. We do not observe the content of any work.
@@
On average, a worker assigned to receive Copilot used it at least once in 45% of the study weeks. Usage rates varied widely across firms, from a low of 21% of treated study workers using Copilot each week to a high of 75%. This variation is consistent with other surveys of AI adoption across firms (Bonney et al., 2024; McElheran et al., 2024), but the variation is notable given our sample is all large firms that were interested in being early adopters of this tool and participating in a study of its use. Industry and variations in workers’ roles (proxied by pre-study worker behavior) explain some of these differences. Nonetheless, indicators for a worker’s firm remain by far the most powerful predictor of usage, suggesting an important role for firm-specific differences that we are unable to measure such as training or managerial practices.
We find clear evidence that workers who use Copilot are beginning to change their work patterns in meaningful ways. During this early adoption phase, we see larger changes in behaviors that workers can adjust independently and less movement in behaviors that require coordination with colleagues.
@@
Using Copilot regularly caused workers to spend nearly 3 hours less per week working on emails (1.4 hours less in the intent-to-treat estimates). This shift represents a 25% time savings from the 11.5 hours per week that the average studied worker spent on email in the pre-period. Copilot users also shifted when they worked on email, consolidating email work to open up two more hours of extended focus time each week. [^2] Appendix B.1 details how we define each of these outcomes. We find suggestive evidence that workers are completing the documents they author 6-20% faster. In contrast, while workers are most likely to use Copilot during video meetings, we find no changes for Copilot users in the time spent in meetings or the types of meetings attended. Similarly, we find only small and statistically insignificant increases in total time spent working in Word and the number of documents authored, indicating that workers have not shifted the nature of their work or taken on new responsibilities. We find few substantive differences in these patterns among workers with more close coworkers who also have access to Copilot or within firms with higher adoption rates, suggesting that larger shifts in responsibilities require time and broad institutional efforts, not just local team coordination.
@@
This work joins a small set of early studies of generative AI in real workplaces, including Brynjolfsson, Li and Raymond (2025), Cui et al. (2024), Otis et al. (2023), and Toner-Rodgers (2024). These real-world studies complement a larger body of work understanding the impact of generative AI on controlled tasks in a lab setting. [^3] See, for example, Peng et al. (2023), Spatharioti et al. (2023), Vaithilingam, Zhang and Glassman (2022), Campero et al. (2022), Noy and Zhang (2023), Dell’Acqua et al. (2023), Cambon et al. (2023) These earlier studies all focus on the impact of AI tools on specific tasks or narrow categories of workers. Surveys indicate rapid adoption of generative AI at work across a much broader set of workers; Humlum and Vestergaard (2025) find 41% of Danish workers in exposed occupations used ChatGPT at work by the end of 2023 and Blandin, Bick and Deming (2024) find 24% of a representative sample of all U.S. workers used generative AI at work by August 2024. However, these surveys cannot measure the impact of generative AI use on work behavior. This study aims to bridge some of the gap between these two lines of research: measuring both take up and impacts of use for a generative AI tool on a broad set of workers, albeit at the cost of clear measures of worker productivity.
1 Patterns of Copilot Adoption
@@
In this study, firms gave treated workers access to Copilot, but did not require them to use it in specific work tasks. [^4] Additional details on the experiment are in Appendix A. The median treated worker used Copilot in 42% of the weeks we followed them (the mean was 45%). [^5] Note that the data reflect engagement with Copilot, not with generative AI tools overall. These adoption rates may not be indicative of broader or long-run trends. During this early adoption period workers had less familiarity with the capacities of generative AI and the tools themselves were less developed, which may have lowered usage. However, firms may have recruited workers for this study who were particularly well suited to being early and enthusiastic adopters. Figure 1(a) shows the distribution of these individual usage rates. 2% of treated workers used Copilot every week, while 6% never used it at all. Figure 1(b) plots the share of treated workers using Copilot for each week since their firm allocated the experimental licenses. Usage trailed off modestly after a peak of 60% when workers were first granted access, presumably due to a combination of people wanting to try the new tool and the introductory training sessions offered by some firms. In the final two months, overall usage held steady at around 40%, for a mean usage rate of 44% of workers across the study.666 The mean weeks of Copilot use across workers differs slightly from the mean share of workers using Copilot in each event week because not all workers were followed for exactly six months. See Appendix A for more details on the experimental design.
@@
Figure 1(b) also shows usage rates over time separately for Microsoft 365 Copilot usage in three programs: Teams, Outlook, and Word. Copilot can be accessed through Outlook, Teams, Word, Excel, PowerPoint, and as a separate M365 Chat, and the ”any” line in Figure 1(b) reflects total use across all apps. We focus our impact analysis on three applications because the early phase of prompt-based generative AI was most applicable to the tasks in these tools. Workers were most likely to use the Copilot in Teams, which can capture transcripts of video meetings and analyze them to generate meeting summaries and follow-up items or answer user queries like “what was discussed in the last 10 minutes?”. The next most common use was through Outlook, where Copilot can summarize email chains, review a worker’s inbox and flag emails that require a response, answer questions about information contained in emails more flexibly than a keyword search, and generate draft emails in response to prompts. The Word Copilot can draft text from prompts and answer questions about the contents of documents. This Copilot had the lowest usage of the three we study, partially because not all studied workers used Word regularly in their jobs (see Appendix B).
@@
Note: Figures (a) and (c) calculate the proportion of weeks in the post-period that each worker used Copilot and show the distribution across (a) workers overall or (c) workers by industry. In Figure (c), each solid dot represents 5 workers; lighter dots represent fewer than 5 workers. Mean industry usage rates are highlighted in red. Figure (b) shows average usage rates across all treated users over time. Figure (d) is a binscatter for the average share of Copilot users in a specific app in the average week grouped by the overall Copilot usage rate of the firm.
Some of the individual heterogeneity is also reflected in average firm usage rates, which ranged from as low as 21% to as high as 75%. [^7] Appendix Figure 3(a) shows how usage varied over time across firms. We do not observe many potentially important sources that likely drive this firm variation. For example, some firms likely provided training to workers on Copilot useage or may have altered managerial practices. However, we can look at how much of the variation in adoption is explained by (the proxies we have for) the type of work done at different firms (firm industry), the type of work done by study participants (pre-experiment individual work patterns [^8] Weekly averages of: emails read, unique email threads replied to, Teams meeting time, Teams meetings attended, time in Word, documents edited or read before completion, number of unique email recipients, number of unique people met 1-on-1.), and how firms distributed their Copilot licenses (the share of a worker’s coworkers who have a license). We cannot estimate the causal effect of any of these factors, only the extent to which they explain the observed variation.
@@
Share of coworkers 0.044 0.053 0.099 0.038
with Copilot (0.024) (0.024) (0.023) (0.024)
@@
Firm FEs Yes Yes Yes
Industry FEs Yes Yes
@@
Note: This table reports OLS regressions where the dependent variable is the share of post-period weeks that an individual used Copilot. Coefficients for individual pre-experiment behavior variables are in Appendix Table A1.
@@
Table 1 shows the $R$-squared from regressions of each treated worker’s average Copilot usage rate over the treatment period on various individual and firm metrics. The associated coefficients are in Appendix Table A1. Individual pre-experiment activity in Outlook, Teams, and Word explains only 9.3% of the variation. The share of a worker’s coworkers that have Copilot has a statistically significant impact on usage, but the effect is small and adds little explanatory power. Moving from no coworkers to the median of 12.5% would increase a person’s predicted usage by 0.55 percentage points. Firm fixed effects matter more, explaining 18.4% of the variation on their own and raising the $R$-squared from 0.094 to 0.235 when added to the other variables. Adoption rates vary meaningfully across industries (ranging from 30% in Construction and Manufacturing to 54% in Consumer Goods as shown in Figure 1(c)). Nevertheless, industry fixed effects explain less than a third as much of the variation in individual usage as firm fixed effects. [^9] Some of the firm variation could also come from time-of-year effects since firms joined the experiment at different times. In the Appendix we repeat this analysis with user-week level data where we can control for calendar time. The $R$-squared’s are overall much lower, since we are not averaging within a user, but firm fixed effects still add about twice as much explanatory power as individual pre-experiment work patterns.
Given the importance of the firm in determining Copilot usage, we also consider whether workers at firms with high uptake rates overall are also using Copilot differently. Figure 1(d) plots the firm average use intensity of Copilot within each of the three applications we study against the overall firm average Copilot usage. A relationship that parallels the 45°$\degree$ line, as is the case for Teams, indicates that this is an equally popular way to use Copilot at all firms. A flatter relationship, which we see in Outlook and Word, indicates that these applications are relatively more popular at firms with lower overall usage: the rate of Outlook and Word Copilot use rises more slowly than the increase in Copilot use overall. These patterns are consistent with our findings on the impacts of Copilot use. In Outlook and Word, Copilot can assist individual behaviors, so workers can find it useful even in environments where few colleagues are adopting this new tool. At firms with more active users, the additional usage seems to mainly come from the other apps. [^10] Appendix Figure 3(b) shows a similar pattern based on the average number of actions. As Copilot usage at a firm increases, the average number of Teams-Copilot actions among Copilot users increases, but the number of Outlook-Copilot actions is flat.

### v2

@@ Abstract
We present evidence on how generative AI changes the work patterns of knowledge workers using data from a 6-month-long, cross-industry, randomized field experiment. Half of the 7,137 workers in the study received access to a generative AI tool integrated into the applications they already used for emails, document creation, and meetings. We find that access to the AI tool during the first year of its release primarily impacted behaviors that workers could change independently and not behaviors that require coordination to change: workers who used the tool in more than half of the sample weeks spent 3.6 fewer hours, or 31% less time on email each week (intent to treat estimate is 1.3 hours) and completed documents moderately faster, but did not significantly change time spent in meetings.
@@
Recent advances in generative AI have opened new possibilities for using this technology to assist with or automate a variety of tasks. Early studies have already shown that these AI can substantially increase workers’ productivity in targeted tasks. [^1] See, e.g., Brynjolfsson, Li and Raymond (2025), Peng et al. (2023), Noy and Zhang (2023), Dell’Acqua et al. (2023). Over time, these advances have the potential to shift the nature of jobs (Eloundou et al., 2024) and perhaps the macroeconomy (Autor (2024), Acemoglu (2025)). However, the transition from the creation of new technologies to effective, widespread adoption in the workplace takes time (Brynjolfsson, Rock and Syverson, 2021). In lab experiments, participants are asked to focus on completing a single task and directed to use AI tools that are well-suited to that task. Outside the lab, workers must learn on their own which parts of their job benefit from the use of new tools and gradually change their habits of work. Larger productivity gains often require organizational changes to redesign processes and shift responsibilities across workers (Bresnahan, Brynjolfsson and Hitt, 2002).
@@
In this paper, we present early evidence on how generative AI tools are changing the way knowledge workers do their jobs. Between September 2023 and October 2024, Microsoft partnered with 66 large firms to run a cross-industry field experiment to measure how access to an integrated generative AI tool, Microsoft 365 Copilot (hereafter, Copilot), changed work patterns. Each firm recruited workers for the study; workers were then randomly assigned to either receive access to this new generative AI tool or continue working with their current technologies. Participants in the study span many occupations, but all participating workers do some kind of work that involves heavy use of the Microsoft Office tools into which Copilot was integrated (e.g., emailing, video meetings, chat, document creation). Firms were asked to maintain the random assignment for six months, giving treated workers sufficient time to explore the new tool and integrate it into their daily work.
@@
The study took place during the early rollout of Copilot, when firms were restricted to purchasing a small number of licenses. [^2] Each company was able to purchase 300 licenses. In all cases, this represents only a small fraction of their total employees who used Microsoft’s Office suite. This particular generative AI tool was new for all workers, and the median worker has only one close coworker with access to the tool (11.9% of their team). Treated workers were free to use or not use the tool during their regular work. Treated and control workers may have also used other generative AI tools during the study (for example, ChatGPT had been publicly available for nearly a year by the beginning of the study), but access to Copilot could make it easier for workers to use generative AI at work. Copilot integrates AI into applications these workers were already using regularly in their jobs, allowing some new uses, like summarizing emails or searching for information within work content, while potentially reducing the effort cost of using generative AI to create content within documents and emails.
@@
We use highly detailed telemetry data collected by Microsoft products to track the behavior of the over 7,000 studied workers during the experiment and for several months before the introduction of the new tool. These telemetry measures capture the actions of workers, including time spent reading emails, attending video meetings, and time from creation to completion of documents. We do not observe the content of any work.
@@
Over 90% of the workers assigned to receive Copilot used it at least once in the 6 month period after treatment, while usage intensity varied widely. The average treated worker used Copilot at least once in 41% of the study weeks, with a standard deviation of 30%. Usage rates varied widely across firms. In the lowest-use firm, treated workers averaged using Copilot in only 6.3% of post-period weeks. In the highest-use firm, treated workers averaged using Copilot in 75% of the post-period weeks. This variation is consistent with other surveys of AI adoption across firms (Bonney et al., 2024; McElheran et al., 2024), but the variation is notable given our sample is all large firms that were interested in being early adopters of this tool and participating in a study of its use. Industry differences and variations in workers’ roles (proxied by pre-study worker behavior) explain some of the heterogeneous usage. Nonetheless, indicators for a worker’s firm remain by far the most powerful predictor of usage, suggesting an important role for firm-specific differences that we are unable to measure, such as training or managerial practices.
@@
We find clear evidence that workers who use Copilot begin to change their work patterns in meaningful ways. During this early adoption phase, we see larger changes in behaviors that workers can adjust independently and less movement in behaviors that require coordination with colleagues. Using Copilot in at least half the post-period weeks caused workers to spend over 3.5 fewer hours per week working on emails (1.3 fewer hours in the intent-to-treat estimates). This shift represents a 31% time savings from the 11.7 hours per week that the average studied worker spent on email in the pre-period. Copilot users also shifted when they worked on email, consolidating email work into blocks that then open up four more hours of extended focus time each week. [^3] Appendix B.1 details how we define each of these outcomes. We find suggestive evidence that treated workers who adopt the tool complete the documents they author 5-25% faster. In contrast, while workers are most likely to use Copilot during video meetings, we find no changes for Copilot users in the time spent in meetings and little shift in the types of meetings attended. Similarly, we find only small and statistically insignificant increases in total time spent working in Word and the number of documents authored, indicating that workers did not shift the nature of their work or take on new responsibilities. We find few substantive differences in these patterns among workers with more close coworkers who also had access to Copilot, suggesting that larger shifts in responsibilities require time and broad institutional efforts, not just local team coordination.
@@
This work joins a small set of early studies of generative AI in real workplaces, including Brynjolfsson, Li and Raymond (2025), Cui et al. (2024), Otis et al. (2023), and Toner-Rodgers (2024). These real-world studies complement a larger body of work understanding the impact of generative AI on controlled tasks in a lab setting. [^4] See, for example, Peng et al. (2023), Spatharioti et al. (2023), Vaithilingam, Zhang and Glassman (2022), Campero et al. (2022), Noy and Zhang (2023), Dell’Acqua et al. (2023), Cambon et al. (2023). These earlier studies all focus on the impact of AI tools on specific tasks or narrow categories of workers. Surveys indicate rapid adoption of generative AI at work across a much broader set of workers, but cannot measure worker behavior; Humlum and Vestergaard (2025a) find 41% of Danish workers in exposed occupations used ChatGPT at work by the end of 2023 and Blandin, Bick and Deming (2024) find 24% of a representative sample of all U.S. workers used generative AI at work by August 2024. In follow up work, Humlum and Vestergaard (2025b) find that this widespread use of AI chatbots at work has not had any measurable impact on worker earnings or hours of work, consistent with our mixed results on worker behavior. This study aims to bridge some of the gap between these two lines of research: measuring both take up and impacts of use for a generative AI tool on a broad set of workers, albeit at the cost of clear measures of worker productivity. Our findings on individual time use clarify why workers might be eager to use these tools at work even without large impacts on earnings.
1 Treatment Assignment and Patterns of Copilot Adoption
At the time the experiment was run, firms could purchase at most 300 M365 Copilot licenses. The Microsoft marketing team recruited firms to join the experiment with the understanding that at least 50 of their licenses would be allocated at random between workers. Firms were asked to recruit at least twice the number of workers for the study as the number of Copilot licenses that would be randomly assigned. Firms gave treated workers access to Copilot, but did not require them to use it in specific work tasks. [^5] Additional details on the experiment are in Appendix A.
@@ 1 Treatment Assignment and Patterns of Copilot Adoption
The median treated worker used Copilot in 39% of the weeks we followed them (the mean was 41%). [^6] Note that the data reflect engagement with Copilot, not with generative AI tools overall. These adoption rates may not be indicative of broader or long-run trends. During this early adoption period workers likely had less familiarity with the capacities of generative AI and the tools themselves were less developed, which may have lowered usage. However, firms may have recruited workers for this study who were particularly well suited to being early and enthusiastic adopters. Figure 1(a) shows the distribution of these individual usage rates. 1.6% of treated workers used Copilot every week, while 10% never used it at all. Figure 1(b) plots the share of treated workers using Copilot for each week since their firm allocated the experimental licenses. Usage trailed off modestly after a peak of 55% when workers were first granted access, presumably due to a combination of people wanting to try the new tool and the introductory training sessions offered by some firms. In the final two months, overall weekly usage held steady at around 37% of workers.
@@
Figure 1(b) also shows usage rates over time separately for Microsoft 365 Copilot usage in three programs: Teams, Outlook, and Word. Copilot can be accessed through Outlook, Teams, Word, Excel, PowerPoint, and as a separate M365 Chat, and the “any” line in Figure 1(b) reflects total use across all apps. We focus our impact analysis on three applications because the early phase of prompt-based generative AI was most applicable to the tasks in these tools. Workers were most likely to use Copilot in Teams, which can capture transcripts of video meetings and analyze them to generate meeting summaries and follow-up items or answer user queries like “what was discussed in the last 10 minutes?”. The next most common use was through Outlook, where Copilot can summarize email chains, review a worker’s inbox and flag emails that require a response, answer questions about information contained in emails more flexibly than a keyword search, and generate draft emails in response to prompts. The Word Copilot can draft text from prompts and answer questions about the contents of documents. This Copilot had the lowest usage of the three we study, partially because not all studied workers used Word regularly in their jobs (see Appendix B). Copilots for Powerpoint and Excel had lower usage rates at the time of the study.
@@
Note: Figures (a) and (c) show the proportion of weeks in the post-period that each worker used Copilot and show the distribution across (a) workers overall or (c) workers by industry. In Figure (c), each solid dot represents 8 workers; lighter dots represent fewer than 8 workers. Mean industry usage rates are highlighted in red. Figure (b) shows average usage rates across all treated users over time. Figure (d) is a binscatter plotting the mean proportion of firms’ Copilot users who are using Copilot within Outlook, Teams, and Word against firms’ overall average weekly Copilot use.
Some of the individual heterogeneity is also reflected in average firm usage rates, which ranged from as low as 6.3% to as high as 75%. [^7] Appendix Figure 3(a) shows how usage varied over time across firms. We do not observe many potentially important sources that likely drive this firm variation. For example, some firms likely provided training to workers on Copilot usage or may have altered managerial practices. However, we can look at how much of the variation in adoption is explained by (the proxies we have for) the type of work done at different firms (firm industry), the type of work done by study participants (pre-experiment individual work patterns [^8] Weekly averages of: emails read, unique email threads replied to, Teams meeting time, Teams meetings attended, time in Word, documents edited or read before completion, number of unique email recipients, number of unique people met 1-on-1.), and how firms distributed their Copilot licenses (the share of a worker’s coworkers who have a license). We cannot estimate the causal effect of any of these factors, only the extent to which they explain the observed variation.
@@
Share of coworkers 0.030 0.055** 0.035
with Copilot (0.040) (0.021) (0.037)
@@
Firm FEs Yes Yes
Industry FEs Yes Yes Yes
Start-month FEs Yes
@@
Note: This table reports OLS regressions where the dependent variable is the share of post-period weeks that an individual used Copilot. Because 91% of treated individuals used Copilot at least once, most of the variation is on the intensive margin. The $all\right>$ row indicates the regression has pre-experiment work behavior included, with coefficients reported in Appendix Table A2. Standard errors are clustered by firm.
@@
Table 1 shows the $R$-squared from regressions of each treated worker’s average share of weeks of Copilot usage over the treatment period on various individual and firm metrics. The associated coefficients are in Appendix Table A2. Individual pre-experiment activity in Outlook, Teams, and Word explains only 11.8% of the variation. The share of coworkers that have Copilot adds little explanatory power. [^9] The implied effect on usage is also quite small. The larger .055 coefficient implies that moving from no coworkers to the median of 11.9% would increase a person’s predicted share of Copilot usage weeks by 0.65 percentage points. Firm fixed effects matter more, explaining 25.6% of the variation on their own and raising the $R$-squared from 0.119 to 0.295 when added to the other variables. Average shares of weekly use vary meaningfully across industries (ranging from 27% in Construction and Manufacturing to 48% in Telecommunications as shown in Figure 1(c)). Nevertheless, industry fixed effects explain less than a fifth as much of the variation in individual usage as firm fixed effects. Some of the firm variation could also come from time-of-year effects since firms joined the experiment at different times. While adding fixed effects for the start month explains some variation, it is still substantially less than the firm fixed effects. [^10] In the Appendix we repeat this analysis with user-week level data where we can control for calendar time. The $R$-squared’s are overall much lower, since we are not averaging within a user, but firm fixed effects still add almost twice as much explanatory power as individual pre-experiment work patterns.
Given the importance of the firm in determining Copilot usage, we also consider whether workers at firms with high uptake rates overall are using Copilot in different applications from those in low takeup firms. Figure 1(d) plots the fraction of a firm’s Copilot users who use Copilot within each application (averaged over weeks) against the firm’s overall average Copilot usage. If patterns of use were independent of firm updake, these relationships would be flat: the same share of Copilot users would use the tool through each application regardless of their firms’ overall usage. Instead, we see that workers in firms with low uptake rates are particularly likely to use Copilot through Outlook and less likely to use Copilot through Teams. These patterns suggest that Teams Copilot, and to a lesser extent Word Copilot, may be particularly valuable when multiple workers use them together. Note that while the share of Copilot users using Outlook Copilot shrinks as firm updake increases, the absolute use (the product of overall takeup and application-specific use) still increases.